Pb(NO3)2 + 2 NaI = PbI2 + 2 NaNO

Suppose in the reaction in which 100 grams of lead(II) nitrate were all used up and only 100 grams of lead(II)iodide were obtained in the experiment, what is the percent yield? 

1
Expert's answer
2022-03-08T17:38:04-0500

Pb(NO3)2 + 2 NaI = PbI2 + 2 NaNO3

According to the equation, n (PbI2) = n (Pb(NO3)2).

n = m / M

M (Pb(NO3)2) = 331.2 g/mol

M (PbI2) = 461 g/mol

n (Pb(NO3)2) = m / M = 100 / 331.2 = 0.3 mol

n (PbI2) = n (Pb(NO3)2) = 0.3 mol

m (PbI2)theoretical = n x M = 0.3 x 461 = 138.3 g

% (PbI2) = (100 x 100) / 138.3 = 72.3%